# Beam tracing and profile evolution for localised beams in inhomogeneous plasmas

### v2

@@ III.1 Gaussian profiles
We will thus end up with the same amplitude evolution for $A\alpha^{(0)}$ instead of $\alpha^{(0)}$, using $\Psi_{\mu\nu}+S_{\mu\rho}\Phi_{\rho\sigma}S_{\rho\nu}$ instead of $\Psi_{\mu\nu}$. Thus, the special case of applying a Gaussian beam profile to eq. 23 is consistent with our beam tracing equations.
As a final note, we want to point out that without $A(\tau)$, our Gaussian beam solution is solution with an eigenvalue of $\frac{1}{2}\Phi_{\mu\nu}\partial K_{\mu}\partial K_{\nu}H$. We are also free to choose any symmetric, complex $\bm{\Phi}$, as the requirement that our Gaussian profile localises our beam only applies to $\bm{\Psi}_{\mathrm{tot}}$. This detail will be important after we construct our ladder operators. In particular, note that if we define:
@@
then we are assured that $\bm{\Phi}=\bm{\Psi}_{w,\mathrm{tot}}-\bm{\Psi}_{w}$. This is a neat choice as we do not care about anything other than the components along $\bm{ww}$ in the first place.
III.2 General solution – ladder operators
Noticing that our PDE is similar to the multidimensional Hermite DE, we have to generalise the approach to solving the DE. Our approach is similar to LeenLeen2016, but we manage to construct generalised ladder operators for this system, which work with generic $\tau$-dependent matrix coefficients. This is a necessary feature in our case, as our matrix coefficients most definitely depend on $\tau$.
We first consider an operator of the form $\hat{\mathcal{L}}_{I}=u_{I}^{\mu}(\tau)\partial_{\mu}$. We want this operator to commute with $\hat{\mathcal{D}}$, such that it can be used as a ladder operator to generate eigenfunctions of $\hat{\mathcal{D}}$. We find that the commutator provides us with the following ODE:
$\begin{split}[\hat{\mathcal{D}},\hat{\mathcal{L}}_{I}]&=\frac{\mathrm{d}u_{I}^{\mu}}{\mathrm{d}\tau}\partial_{\mu}-u_{I}^{\mu}T_{\mu\nu}\partial_{\nu}=0,\\ \frac{\mathrm{d}u_{I}^{\nu}}{\mathrm{d}\tau}&=u_{I}^{\mu}T_{w,\mu\nu}.\end{split}$ (35)
This ODE can be expressed in the following vector form:
$\left(\frac{\mathrm{d}\bm{u}_{I,w}}{\mathrm{d}\tau}\right)_{w}=\bm{u}_{I}(\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}\bm{\nabla}_{\bm{K}}H)_{w}.$ (36)
@@ III.2 General solution – ladder operators
$\hat{\mathcal{L}}_{I}(\tau)=\bm{u}_{I}(\tau)\cdot\bm{\nabla}_{w},$ (37)
where:
$\hat{\mathcal{D}},\hat{\mathcal{L}}_{I}]=0.$ (38)
We can presciently construct an adjoint operator, which also happens to be the inverse operator of our original ladder operator. This operator turns out to be something of the form $\hat{\mathcal{R}}_{I}=h_{I}^{\nu}(\tau)\partial_{\nu}+v^{\nu}_{I}(\tau)x_{\nu}$. We find that:
$\hat{\mathcal{D}},\hat{\mathcal{R}}_{I}]=\frac{\mathrm{d}v_{I}^{i}}{\mathrm{d}\tau}w_{i}+\frac{\mathrm{d}h_{I}^{i}}{\mathrm{d}\tau}\partial_{i}-\mathrm{i}\partial K_{\mu}\partial K_{\nu}Hv_{I}^{\mu}\partial_{\nu}+w_{\mu}T_{\mu\nu}v_{I}^{\nu}-h_{I}^{\mu}T_{\mu\nu}\partial_{\nu}=0.$ (39)
We can utilise our freedom to define $v^{\nu}_{I}(\tau)$ by:
$\frac{\mathrm{d}v_{I}^{\mu}}{\mathrm{d}\tau}=-T_{w,\mu\nu}v_{I}^{\nu}.$ (40)
If we take $h_{I}^{\mu}(\tau)=M_{\mu\nu}(\tau)v^{\nu}_{I}(\tau)$, we then arrive at another ODE for $M_{\mu\nu}$, which is given in matrix notation by:
$\begin{split}\left(\frac{\mathrm{d}\bm{M}}{\mathrm{d}\tau}\right)_{w}&=\big(\mathrm{i}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\\ &+\bm{M}\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\bm{\Psi}\bm{M}\\ &+\bm{M}\bm{\nabla}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}H\bm{M}\big)_{w},\end{split}$ (41)
where our total derivative includes the changes in the matrix due to the $\tau$-dependent basis. Without loss of generality, if $\bm{M}$ obeys the unprojected equation, this would imply that it is a solution to eq. 41 as well. Specifically, if $\bm{M}$ satisfies:
$\begin{split}\frac{\mathrm{d}\bm{M}}{\mathrm{d}\tau}&=\mathrm{i}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\\ &+\bm{M}\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\bm{\Psi}\bm{M}\\ &+\bm{M}\bm{\nabla}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}H\bm{M},\end{split}$ (42)
then it is a solution to eq. 41. Note that $M_{\mu\nu}$ must be symmetric, and that this bears a strong resemblance to our equation for the evolution of a Gaussian profile, eq. 32. In fact it is almost identical to the equation for the evolution of the inverse profile, given by:
$\begin{split}\frac{\mathrm{d}\bm{\Phi}^{-1}}{\mathrm{d}\tau}&=\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\\ &+\bm{\Phi}^{-1}\bm{\Psi}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\bm{\Psi}\bm{\Phi}^{-1}\\ &+\bm{\Phi}^{-1}\bm{\nabla}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}H\bm{\Phi}^{-1}.\end{split}$ (43)
thus, $\bm{M}=\mathrm{i}\bm{\Phi}^{-1}$ is a valid solution in general, for any $\bm{\Phi}$ where both $\bm{\Psi}+\bm{\Phi}$ and $\bm{\Psi}$ obey the beam tracing equations. Note that $\bm{\Phi}$ need not be full-rank, due to the fact that we only need to satisfy the ODE along the width of the beam. In the case that $\bm{\Phi}$ is not full-rank, $\bm{\Phi}^{-1}$ should be interpreted as the Moore-Penrose inverse. It would not be full rank if we use the prescription in eq. 34, which would be cleanest prescription to use from a computational standpoint.
To summarise, our ladder operators are:
$\begin{split}\hat{\mathcal{L}}_{I}&=\bm{u}_{I}(\tau)\cdot\bm{\nabla}_{w},\\ \hat{\mathcal{R}}_{I}&=\bm{v}_{I}(\tau)\cdot(\bm{w}+\mathrm{i}\bm{\Phi}^{-1}\bm{\nabla}_{w}).\end{split}$ (44)
These two operators commute with our differential operator, allowing us to exactly calculate the evolution of any beam profile matching their boundary conditions. The index ‘I’ does not have to only specify two vectors, in theory we can use an arbitrary number of ladder operators, all with different initial directions to reconstruct any boundary profile one might wish to reconstruct. Doing this in practice might be tricky, but if it can be done, there will be a very significant computational speedup compared to decomposing a profile into Hermite polynomials, as we only have to evolve one solution, instead of a potentially large set of different solutions.
As an interesting note, the stationary states of our ladder operators are $f=C$ for $\hat{\mathcal{L}}_{I}$ and $f=C\exp\left(\frac{\mathrm{i}}{2}\bm{w}\cdot\bm{\Phi}\cdot\bm{w}\right)$ for $\hat{\mathcal{R}}_{I}$. These are precisely the constant profile and Gaussian profile solutions, suggesting that we can use them to generate two families of solutions each. We can also get these families to coincide by matching their boundary conditions. This is similar to how the Hermite polynomials and Hermite functions are related by their respective ladder operators.
@@ III.3 Gauss-Hermite boundary conditions
As mentioned earlier, we would like these boundary profiles to be Hermite polynomials. We can do this simply by making appropriate choices for $\bm{u}_{I}(0)$, $\bm{v}_{I}(0)$, $\bm{\Phi}$ and $\bm{\Psi}$. We are free to choose any arbitrary scaling for our Hermite polynomials due to this freedom, which we will utilise to choose a neat set of boundary conditions in section V.1.2, in this case yielding physicist’s Hermite polynomials instead of probabilist’s Hermite polynomials. However, the subsequent prescription we provide reduces to the ‘elegant’ Hermite polynomialspampaloni2004 at the plasma boundary, which are probabilist’s Hermite polynomials. This will serve as a useful demonstration of our approach, and it serves as a natural basis for decomposing an arbitrary electric field at the boundary.
@@ III.3.1 Case 1
Suppose that we want to generate our Hermite polynomials at the boundary by the following ladder operators:
@@
where $H_{n}(x)H_{m}(y)=\hat{\mathcal{R}}^{n}_{x}(\tau=0)\hat{\mathcal{R}}^{m}_{y}(\tau=0)(1)$, and $\bm{x}(0)=\sqrt{-\mathrm{i}\bm{\Psi}_{w,0}}\bm{w}$. We implicitly assume that $\bm{\Psi}_{w,0}$ is diagonalisable at the boundary such that we can take the square root.
Converting to dimensionless coordinates $\bm{x}(\tau)=\sqrt{\mathrm{i}\bm{\Phi}}\bm{w}$, we have:
@@
where $\sqrt{\bm{\Phi}}$ is defined to be $\sqrt{\bm{\Phi}}\sqrt{\bm{\Phi}}=\bm{\Phi}$. This is not always possible, but given that we are only trying to match to boundary conditions, we take $\bm{\Phi}(0)$ to be a diagonalisable symmetric matrix.
We then set $\bm{\Phi}(\tau=0)=-\bm{\Psi}_{w,0}$ in order to reproduce the ladder operator for the Gauss-Hermite profiles at the boundary. This means that $\bm{\Psi}_{\mathrm{tot}}(\tau=0)=0$, so we thus just need to evolve this alongside $\bm{\Psi}$ to be able to keep track of the evolution of $\bm{\Phi}$. Given the diagonalised basis of $\bm{\Psi}_{w,0}$, call this $\hat{\bm{e}}_{I}$, we then set:
@@
in order for our ladder operators to reduce to that of the normalised Hermite polynomials.
With these choices, the evolution of the profile given by the Hermite polynomial
$H_{n}\left(\sqrt{-\mathrm{i}\Psi_{0,xx}}w_{x}\right)H_{m}\left(\sqrt{-\mathrm{i}\Psi_{0,yy}}w_{y}\right)$ is simply:
$\hat{\mathcal{R}}_{x}^{n}\hat{\mathcal{R}}_{y}^{m}(1),$ (48)
and it reduces to the Hermite polynomials at the boundary.

### v3

@@ III.1 Gaussian profiles
We will thus end up with the same amplitude evolution for $\alpha(\tau)A^{(0)}$ instead of $A^{(0)}$, using $\bm{\Psi}+\bm{\Phi}$ instead of $\bm{\Psi}$. Thus, the special case of applying a Gaussian beam profile to eq. 24 is consistent with our beam tracing equations.
III.2 General solution – ladder operators
In this section, we will present the evolution of a generic starting profile $\mathcal{P}$. Our approach is similar to Leen et. al[Leen2016], but we instead construct generalised ladder operators for this system, which work with generic $\tau$-dependent tensor coefficients. This is a necessary feature in our case, as our tensor coefficients most definitely depend on $\tau$. In a similar fashion to Leen et. al, in subsequent derivations we utilise the capitalised index ‘I’ to denote the ladder operator indexed by ‘I’. Although this is certainly not necessary for the construction of the individual ladder operators, but it is useful book-keeping notation, and makes it explicitly clear that we expect to construct not just a single pair of ladder operators, but sets of ladder operator pairs. To be precise, in n-dimensions, we require one set of n independent ladder operators for $\hat{\mathcal{L}}_{I}$ which generate one family of solutions, and another such set to generate the other family of solutions from $\hat{\mathcal{R}}_{I}$. In the proceeding subsection, we will be outlining the definition and evolution of two ladder operators, $\hat{\mathcal{L}}_{I}$ and $\hat{\mathcal{R}}_{I}$. It is clear from the context when $\hat{\quad}$’ refers to the ladder operator or a unit vector, as $\hat{\mathcal{L}}_{I}$, $\hat{\mathcal{R}}_{I}$ and $\hat{\mathcal{D}}$ will be the main operators used subsequently. By the end of the section, we will have demonstrated that from appropriately chosen stationary states, we can generate families of solutions from either $\hat{\mathcal{L}}_{I}$ or $\hat{\mathcal{R}}_{I}$. In our particular case of n=2, we will construct two ladder operators indexed by ‘I’, that is one pair for $\hat{\mathcal{L}}_{I}$ and one pair for $\hat{\mathcal{R}}_{I}$, where ‘I’ could be ‘x’ or ‘y’.
In this particular ladder operator approach, we are interested in ladder operators that commute with $\hat{\mathcal{D}}$, unlike conventional ladder operator approaches. We will still obtain a ladder of states, but they will no longer be classified hierarchically by their eigenvalues, for example in Leen et. al[Leen2016], as our definition subsumes the exponential eigenvalue modulation into the solutions themselves.
We first construct an operator of the form $\hat{\mathcal{L}}_{I}=u_{I}^{\mu}(\tau)\frac{\partial}{\partial w_{\mu}}$, where $u_{I}^{\mu}$ is an arbitrary vector. We want this operator to commute with $\hat{\mathcal{D}}$, such that it can be used as a ladder operator to generate solutions of $\hat{\mathcal{D}}$. The commutator of $\hat{\mathcal{L}}_{I}$ with $\hat{\mathcal{D}}$ is given by:
$\hat{\mathcal{D}},\hat{\mathcal{L}}_{I}]=\frac{\mathrm{d}u_{I}^{\mu}}{\mathrm{d}\tau}\frac{\partial}{\partial w_{\mu}}-u_{I}^{\mu}T_{w,\mu\nu}\frac{\partial}{\partial w_{\nu}}.$ (36)
In order for the commutator to evaluate to 0, our ladder operator components must satisfy the following ODE:
$\frac{\mathrm{d}u_{I}^{\nu}}{\mathrm{d}\tau}=u_{I}^{\mu}T_{w,\mu\nu}.$ (37)
Our choice of $\bm{u}_{I}(0)$ is still ultimately completely arbitrary, and depend on the specific boundary conditions we are interested in. After solving for $\bm{u}_{I}(\tau)$, our ladder operator is given by:
$\hat{\mathcal{L}}_{I}(\tau)=\bm{u}_{I}(\tau)\cdot\bm{\nabla}_{w}.$ (38)
We can presciently construct another operator that will turn out to operate similar to an inverse operator. This operator turns out to be something of the form $\hat{\mathcal{R}}_{I}=h_{I}^{\nu}(\tau)\frac{\partial}{\partial w_{\nu}}+v^{\nu}_{I}(\tau)w_{\nu}$, where both $v_{I}^{\nu}$ and $h_{I}^{\nu}$ are arbitrary vectors. We find that:
$\hat{\mathcal{D}},\hat{\mathcal{R}}_{I}]=\frac{\mathrm{d}v_{I}^{\mu}}{\mathrm{d}\tau}w_{\mu}+\frac{\mathrm{d}h_{I}^{\mu}}{\mathrm{d}\tau}\frac{\partial}{\partial w_{\mu}}-\mathrm{i}\frac{\partial}{\partial K_{\mu}}\frac{\partial}{\partial K_{\nu}}Hv_{I}^{\mu}\frac{\partial}{\partial w_{\nu}}+w_{\mu}T_{\mu\nu}v_{I}^{\nu}-h_{I}^{\mu}T_{\mu\nu}\frac{\partial}{\partial w_{\nu}}.$ (39)
Comparing the coefficients of $\frac{\partial}{\partial w_{\nu}}$ and $w_{\nu}$, we require each set of coefficients to be 0. We thus find that $v^{\nu}_{I}(\tau)$ must be given by:
$\frac{\mathrm{d}v_{I}^{\mu}}{\mathrm{d}\tau}=-T_{w,\mu\nu}v_{I}^{\nu}.$ (40)
We also get an ODE for the evolution of $h^{\nu}_{I}(\tau)$:
$\frac{\mathrm{d}h_{I}^{\mu}}{\mathrm{d}\tau}=h_{I}^{\nu}T_{w,\nu\mu}+\mathrm{i}v_{I}^{\nu}\frac{\partial}{\partial K_{\nu}}\frac{\partial}{\partial K_{\mu}}H.$ (41)
The subsequent mathematics will be much more illuminating if we take $h_{I}^{\mu}(\tau)=M_{\mu\nu}(\tau)v^{\nu}_{I}(\tau)$, as this will expose a deep link to the Gaussian profile. Using eqs. 40 & 41, we then arrive at another ODE for $M_{\mu\nu}$:
$\left(\frac{\mathrm{d}M_{\mu\nu}}{\mathrm{d}\tau}\right)_{w}=\left(\mathrm{i}\frac{\partial}{\partial K_{\nu}}\frac{\partial}{\partial K_{\mu}}H+M_{\rho\nu}T_{w,\rho\mu}+M_{\mu\rho}T_{w,\rho\nu}\right)_{w}$ (42)
[…]
$\begin{split}\left(\frac{\mathrm{d}\bm{M}}{\mathrm{d}\tau}\right)_{w}&=\big(\mathrm{i}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\\ &+\bm{M}\cdot\bm{\Psi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Psi}\cdot\bm{M}\\ &+\bm{M}\cdot\bm{\nabla}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}H\cdot\bm{M}\big)_{w},\end{split}$ (43)
Without loss of generality, if $\bm{M}$ obeys the unprojected equation, this would imply that it is a solution to eq. 43 as well. Specifically, if $\bm{M}$ satisfies:
$\begin{split}\frac{\mathrm{d}\bm{M}}{\mathrm{d}\tau}&=\mathrm{i}\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\\ &+\bm{M}\cdot\bm{\Psi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Psi}\cdot\bm{M}\\ &+\bm{M}\cdot\bm{\nabla}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}H\cdot\bm{M},\end{split}$ (44)
then it is a solution to eq. 43. Note that this bears a strong resemblance to our equation for the evolution of a Gaussian profile, eq. 30. In fact it is almost identical to the equation for the evolution of the inverse profile, given by:
$\begin{split}\frac{\mathrm{d}\bm{\Phi}^{-1}}{\mathrm{d}\tau}&=\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\\ &+\bm{\Phi}^{-1}\cdot\bm{\Psi}\cdot\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}_{\bm{K}}H\cdot\bm{\Psi}\cdot\bm{\Phi}^{-1}\\ &+\bm{\Phi}^{-1}\cdot\bm{\nabla}\bm{\nabla}_{\bm{K}}H+\bm{\nabla}_{\bm{K}}\bm{\nabla}H\cdot\bm{\Phi}^{-1}.\end{split}$ (45)
thus, $\bm{M}=\mathrm{i}\bm{\Phi}^{-1}$ is a valid solution in general, for any $\bm{\Phi}$ where both $\bm{\Psi}+\bm{\Phi}$ and $\bm{\Psi}$ obey the beam tracing equations. Note that $\bm{\Phi}$ need not be full-rank, due to the fact that we only need to satisfy the ODE for the components along the width of the beam. We get no further information from the components that do not lie along the width of the beam at any rate. In the case that $\bm{\Phi}$ is not full-rank, $\bm{\Phi}^{-1}$ should be interpreted as the Moore-Penrose inverse. It would not be full rank if we use the prescription in eq. 47, which would be cleanest prescription to use from a computational standpoint.
To summarise, our ladder operators are:
$\begin{split}\hat{\mathcal{L}}_{I}&=\bm{u}_{I}(\tau)\cdot\bm{\nabla}_{w},\\ \hat{\mathcal{R}}_{I}&=\bm{v}_{I}(\tau)\cdot(\bm{w}+\mathrm{i}\bm{\Phi}^{-1}\cdot\bm{\nabla}_{w}).\end{split}$ (46)
[…]
As a final note, we want to point out that we are free to choose any convenient symmetric $\bm{\Psi}_{\mathrm{tot}}$ for obtaining $\bm{\Phi}$, as the only thing we have set in stone is $\bm{\Psi}$ when we defined our beam envelope. The initial choice of $\bm{\Phi}$ simply affects the initial definition for $\bm{\Psi}_{\mathrm{tot}}$ and vice versa. In particular, note that if we make the following definition:
@@ III.2 General solution – ladder operators
then we are assured that $\bm{\Phi}=\bm{\Psi}_{w,\mathrm{tot}}-\bm{\Psi}_{w}$. This is a neat choice as we do not care about anything other than the components along $\bm{ww}$ in the first place, and will henceforth be the choice of $\bm{\Phi}$ we use in the rest of the manuscript.
Firstly, we have to identify the stationary states of our ladder operators, so we look for the two functions for which $\hat{\mathcal{L}}_{I}f=0$ and $\hat{\mathcal{R}}_{I}g=0$. On top of this, the solutions need to be solutions to $\mathcal{D}$. [^4] Notice that if we had a function f satisfying $\hat{\mathcal{L}}_{I}f=0$, then it is clear that $\hat{\mathcal{D}}\hat{\mathcal{L}}_{I}f=\hat{\mathcal{L}}_{I}\hat{\mathcal{D}}f=0$. It is easy to see that $\hat{\mathcal{L}}_{I}$ and $\hat{\mathcal{R}}_{I}$ only have one-dimensional kernels, thus this tells us that $\hat{\mathcal{D}}f=C(\tau)f$. Hence, stationary states of our ladder operators are almost solutions to $\hat{\mathcal{D}}$ in general, and can be made into solutions by redefining to $f^{\prime}=f\exp\left(\int_{0}^{\tau}C(\tau^{\prime})\mathrm{d}\tau^{\prime}\right)$, such that $\hat{\mathcal{D}}f^{\prime}=0$. The same argument trivially follows for $\hat{\mathcal{R}}_{I}$. These are $f=C$ for $\hat{\mathcal{L}}_{I}$ and $g=\alpha(\tau)\exp\left(({\mathrm{i}}/{2})\bm{w}\cdot\bm{\Phi}\cdot\bm{w}\right)$ for $\hat{\mathcal{R}}_{I}$, which are precisely the constant profile and Gaussian profile solutions we were already aware of, suggesting that we can use them to generate a family of solutions to $\hat{\mathcal{D}}$ each. We generate each family of solutions by either repeated application of various $\hat{\mathcal{R}}_{I}$ to the constant profile solution, or repeated application of various $\hat{\mathcal{L}}_{I}$ to the Gaussian profile solution. In particular, we can generate each family of solutions, $f_{nm}$ and $g_{nm}$, in our two-dimensional case(see section III.3 for details) using:
@@
where we choose $\alpha(0)=1$.
To summarise, in this section we have found ladder operator solutions to $\hat{\mathcal{D}}$. In the subsequent subsection, we will utilise the ladder operators to generate solutions which reduce to the Hermite polynomials at the plasma boundary, $\tau=0$. The Hermite polynomial profile is a useful boundary profile as it allows us to easily decompose a complicated electric field into various Gauss-Hermite modes by utilising the fact that the Gauss-Hermite modes form an orthogonal basis. In theory, this allows us to model the evolution of an arbitrary profile by first decomposing it into its Gauss-Hermite mode constituents, and then propagating each mode.
@@ III.3 Gauss-Hermite boundary conditions
As mentioned earlier, we would like these profiles to reduce to Hermite polynomials at $\tau=0$. We can do this simply by making appropriate choices for $\bm{u}_{I}(0)$, $\bm{v}_{I}(0)$, $\bm{\Phi}$ and $\bm{\Psi}$. We are free to choose any arbitrary scaling for our Hermite polynomials due to this freedom, which we will utilise to choose a neat set of boundary conditions in section V.0.2, in this case yielding physicist’s Hermite polynomials instead of probabilist’s Hermite polynomials. However, the subsequent prescription we provide reduces to the ‘elegant’ Hermite polynomials[pampaloni2004] at the plasma boundary, which are probabilist’s Hermite polynomials. This will serve as a useful demonstration of our approach, and it serves as a natural basis for decomposing an arbitrary electric field at the boundary. To do this, we will first express the elegant Hermite polynomials(or any Hermite polynomial convention of one’s choosing) in terms of ladder operators, then match the ladder operators we developed.
@@ III.3.1 Case 1
Suppose that we want to generate probabilist’s Hermite polynomials of a dimensionless variable $\bm{x}$ at the boundary by the following ladder operators, where $\hat{\bm{e}}_{i}$ is a unit vector in a given coordinate basis and $\bm{\nabla}_{x}$ is the gradient operator in $\bm{x}$:
@@
The n-th order probabilist’s Hermite polynomial in the coordinate x is given by $H_{n}(x)=\hat{A}_{x}^{+,n}(1)$, referring to application of $\hat{A}_{i}^{+}$ n times on unity. In 2D, with two independent coordinates x and y, the $(n,m)$-th order Hermite polynomial is unambiguously given by $H_{n}(x)H_{m}(y)=\hat{A}_{x}^{+,n}\hat{A}_{y}^{+,m}(1)$, where we can apply the ladder operators in any order due to the fact that the ladder operators in orthogonal directions commute(see section B for details).
We first convert the ladder operators $\hat{\mathcal{L}}_{I}$ and $\hat{\mathcal{R}}_{I}$ to dimensionless coordinates $\bm{x}(\tau)=\sqrt{\mathrm{i}\bm{\Phi}}\cdot\bm{w}$, in order to directly compare them with the aforementioned ladder operators, $\hat{A}_{x}^{-}$ and $\hat{A}_{x}^{+}$ respectively. We then have:
@@
[…]
in order for our ladder operators to reduce to that of the normalised elegant Hermite polynomials. With these choices, the solution given by:
$f_{nm}=\hat{\mathcal{R}}_{x}^{n}\hat{\mathcal{R}}_{y}^{m}(1),$ (52)
satisfies the boundary condition $f_{nm}(\tau=0)=H_{n}\left(\sqrt{-\mathrm{i}\Psi_{0,xx}}w_{x}\right)H_{m}\left(\sqrt{-\mathrm{i}\Psi_{0,yy}}w_{y}\right)$.